OpenStudy (anonymous):

Can someone help me find the point slope form of the equation For a line that passes through (2,3) and (5,4)

OpenStudy (anonymous):

No one usually helps me on here :( but please? Somebody?

OpenStudy (anonymous):

I can help

OpenStudy (anonymous):

Thanks that would be great :)

OpenStudy (anonymous):

This is your slope formula: \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] When you input your points which are (2,3) and (5,4) [2 being x1, 3 being y1, 5 being x2 and y2 being 4] You will get a slope of: \[\frac{ 1 }{ 3 }\] When finding point slope formula you use the formula: \[y-y _{1}=m(x-x _{1})\] m= slope (You can use the y and x of any ordered pair but if u choose one the other one has to be from the same ordered pair. I am going to use the ordered pair (2,3)) My x1 in this ordered pair is 3 and my y1 is 2. My slope is 1/3 Since I have all my variables, I subsitute which gives me \[y-3=\frac{ 1 }{ 3 }(x-2)\] And I solve My final answer is\[y=\frac{ 1 }{ 3 }x+\frac{ 7 }{ 3 }\]